Find the variation constant, and write a formula that expresses the indicated variation.
n varies directly as q, and n  39 when q  3. Find n
when q  8.
same here, can't quite get the hang of these.

Answer by stanbon(75887)

Find the variation constant, and write a formula that expresses the indicated variation.
n varies directly as q, and n  39 when q  3. Find n
when q  8.
-------------------
n = kq
Find "k":
39 = k*5
k = (5/39)
-----------------
So the equation is n = (5/39)q
----------------
Find n when q=8
n = (5/39)8
n = 40/39
-------------
Cheers,
Stan H.